OpenStudy (anonymous):

How do I find the double integral by using transformation equations. Given x-y =u x+y =v evaluate ſſR (x-y)^4*e^(x+y)dxdy where R is the Square with vertices (1,2), (0,1), (2,1) and (1,0),.... anybody there to help??? ſ ſ is the double integral

OpenStudy (amistre64):

looks like we can start with a simple replacement: (x-y)^4 e^(x+y) --> u^4 e^v swapping the dxdy part tho; we need to define dxdy in terms of dudv, or dvdu

OpenStudy (amistre64):

u = x-y du = 1dx v = x+y dv = 1dy im sure that off

OpenStudy (anonymous):

You need to multiply by the jacobian of the transformation \[dxdy=\frac{\partial(x,y)}{\partial(u,v)}dudv\]

OpenStudy (amistre64):

yes, the jacobian seems to apply here :)

OpenStudy (anonymous):

but what is he limit R need to apply here ??

OpenStudy (anonymous):

R is actually a square vertices (1,2), (0,1), (2,1) and (1,0) ?

OpenStudy (amistre64):

Given x-y =u, and x+y =v and a lot of points (x,y), you can determine the transformation from xy to uv

OpenStudy (amistre64):

(1,2), (0,1), (2,1) and (1,0) x-y =u, and x+y =v 1 2 1 2 (1,2) -> (-1,3) ====================== x-y =u, and x+y =v 0 1 0 1 (0,1) -> (-1,1) etc